OpenStudy (dtan5457):

Trig question check If csc=3/2 and pi/2

OpenStudy (dtan5457):

Sin=2/3 use the pythagorean identity (2/3)^2+cos^2=1 cos^2=1-(2/3)^2 cos^2=5/9 cos=sqrt of 5/3 tan=sin/cos tan=

OpenStudy (dtan5457):

\[-\frac{ 2\sqrt{5} }{ 5 }\]

OpenStudy (dtan5457):

@bibby @Directrix

OpenStudy (dtan5457):

@Loser66 @SolomonZelman

OpenStudy (dtan5457):

i get most of it but my book says the answer is POSITIVE..

OpenStudy (dtan5457):

i would think the domain here is the 2nd quadrant

OpenStudy (misty1212):



OpenStudy (misty1212):

\[\tan(\theta)=-\frac{2}{\sqrt5}\]

OpenStudy (misty1212):

the minus because you are in quad 2

OpenStudy (dtan5457):

yep. knew something was wrong with my book, thanks